Imagine a population that has 179 A1A1 individuals, 184 A1A2 individuals, and 26 A2A2 individuals. Male A1A1 is the preferred mate of females and gets all the matings. What would the genotype frequencies be for the A1A2 offspring of this population?
Answer #1

Answer:

Each individual has two alleles.

179A1A1 = 179*2 = 358 A1 alleles

184 A1A2 = A1 184 & A2 184 alleles

26 A2A2 = 52 A2 alleles

Total alleles = 778

Total A1 alleles = 358 + 184 = 542

Frequency of A1 allele = 542/778 = 0.697

Total A2 alleles = 52 + 184 = 236

Frequency of A2 allele = 236/778 = 0.303

Frequency of A1A2 genotype = 2A1A2 = 2 * 0.697 * 0.303 = 0.422 or 42%